Problem 14-9 Bonita Cosmetics Co. purchased machinery on December 31, 2016, paying $52,200 down and agreeing to pay the balance in four equal installments of $57,600 payable each December 31, An assumed interest of 10% is implicit in the purchase price Prepare the journal entries that would be recorded for the purchase and for (1) the payments and (2) interest on the following dates. (Round answers to O decimal places, e.g. 38,548. If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.) (a) December 31, 2016. (b) December 31, 2017 (c) December 31, 2018 (d) December 31, 2019 (e) December 31, 2020 No. Account Titles and Explanation Debit Credit
Answer #1
No. Date Account titles and Explanation Debit Credit
(a) December 31, 2016 Machinery $ 2,34,784
Cash $       52,200
Loan payable $   1,82,584
(To record purchase of machine)
(b) December 31, 2017 Loan Payable $      39,342
Interest Expense $      18,258
Cash $       57,600
(To record interest payment and loan repayment)
(c) December 31, 2018 Loan Payable $      43,276
Interest Expense $      14,324
Cash $       57,600
(To record interest payment and loan repayment)
(d) December 31, 2019 Loan Payable $      47,603
Interest Expense $        9,997
Cash $       57,600
(To record interest payment and loan repayment)
(e) December 31, 2020 Loan Payable $      52,364
Interest Expense $        5,236
Cash $       57,600
(To record interest payment and loan repayment)
Wotking:
Present Value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.10)^-4)/0.10 i 10%
= 3.169865446 n 4
Present value of annual installment = Annual Installment*Present value of annuity of 1
= $             57,600 * 3.16986545
= $         1,82,584
Loan amortization table:
Period number Beginning Loan (a) Interest Expense (b)=(a)*10% Annual Installment © reduction in principal (d)=©-(b) Ending Loan Balance (e)=(a)-(d)
1 $       1,82,584 $       18,258.42 $      57,600 $ 39,341.58 $ 1,43,242.67
2 $ 1,43,242.67 $       14,324.27 $      57,600 $ 43,275.73 $     99,966.94
3 $     99,966.94 $         9,996.69 $      57,600 $ 47,603.31 $     52,363.64
4 $     52,363.64 $         5,236.36 $      57,600 $ 52,363.64 $               0.00